Strickler Technology is considering changes in its working capital policies to improve its cash flow cycle. Strickler's sales last year were $4,185,000 (all on credit), and its net profit margin was 4%. Its inventory turnover was 7 times during the year, and its DSO was 34 days. Its annual cost of goods sold was $2,450,000. The firm had fixed assets totaling $725,000. Strickler's payables deferral period is 39 days. Assume 365 days in year for your calculations. Do not round intermediate calculations.

a. Calculate Strickler's cash conversion cycle. Round your answer to two decimal places.

b. Assuming Strickler holds negligible amounts of cash and marketable securities, calculate its total assets turnover. Round your answer to two decimal places.

c. Calculate its ROA. Round your answer to two decimal places.

d. Suppose Strickler's managers believe the annual inventory turnover can be raised to 9 times without affecting sale or profit margins. What would Strickler's cash conversion cycle have been if the inventory turnover had been 9 for the year? Round your answer to two decimal places.

What would Strickler's total assets turnover have been if the inventory turnover had been 9 for the year? Round your answer to two decimal places.

What would Strickler's ROA have been if the inventory turnover had been 9 for the year? Round your answer to two decimal places.
Answer #1
Sales last year (all on credit) $4,185,000 Formulas
net profit margin 4%
Inventory turnover 7 times ITO=Cost of goods sold/average inventory
DSO (Days Sales Outstanding) 34 days DSO=(Accounts receivable*365/net credit sales)
cost of goods sold (COGS)p.a. $2,450,000 COGS=Opening inventory+purchases-closing inventory
Fixed assets $725,000
payables deferral period (DPO) 39 days DPO=Accounts payable*365/COGS
a. Cash conversion cycle CCC=Days Inventory Outstanding+Days Sales outstanding-Days Payable outstanding
Now calculating CCC (please see workings below) CCC=DIO+DSO-DPO
=52.14285+34-39
CCC days rounded off to two decimal places as per instructions. 47.14
Working for average inventory
ITO=Cost of goods sold/average inventory
therefore (7=2,450,000/average inventory)
Solving the above we get average inventory=2,450,000/7 =350,000
Working for DIO
Days Inventory Outstanding (DIO) DIO=Average Inventory*365/COGS
=350,000*365/2,450,000
52.14285714
Answer a. Cash conversion cycle is 47.14 days.
b. Assets Turnover ratio ATO= Net Sales/Average Total Assets
=4,185,000/1,464,835.6164
ATO rounded off to two decimal places as per instructions. 2.86
Working for average total assets
Total assets=Fixed Assets+Current Assets
Current Assets=Inventory+Sales Receivables
Current Assets=350,000+389,835.6164=739,835.6164
So, Total Assets=725,000+739,835.6164=1,464,835.6164
Working for accounts receivable amount
DSO=(Accounts receivable*365/net credit sales)
Therefore 34=AR*365/4,185,000
Solving this, we get AR=4,185,000*34/365
Therefore AR=$389,835.6164
c. Return on Assets (ROA) ROA=Net Income/Average Total Assets
=167,400/1,464,835.6164
ROA rounded off to two decimal places as per instructions. 0.11
Working for Net Income
Net profit margin=4% of sales
Therfore Net Income=4% *4,185,000=$167,400
d. Calculations for CCC, ATO, ROA if ITO is 9 times, and Sales and net profit remain same as above
i.Cash conversion cycle in this case CCC=DIO+DSO-DPO
=40.5556+34-39
r/off to two decimal places 35.56
Working for average inventory
ITO=Cost of goods sold/average inventory
therefore (9=2,450,000/average inventory)
Solving the above we get average inventory=2,450,000/9 =272222.2222
Working for DIO
Days Inventory Outstanding (DIO) DIO=Average Inventory*365/COGS
=272,222.2222*365/2,450,000
40.5556
ii. Assets Turnover ratio ATO= Net Sales/Average Total Assets
=4,185,000/1,387,057.8386
ATO rounded off to two decimal places as per instructions. 3.02
Working for average total assets
Total assets=Fixed Assets+Current Assets
Current Assets=Inventory+Sales Receivables
Current Assets=272,222.2222+389,835.6164=662,057.8386
So, Total Assets=725,000+662,057.8386=1,387,057.8386
iii. Return on Assets (ROA) ROA=Net Income/Average Total Assets
=167,400/1,387,057.8386
ROA rounded off to two decimal places as per instructions. 0.12